OpenStudy (anonymous):

-5x + 7y =16 find the value of y if x =8

OpenStudy (anonymous):

-5(8)+7y=16 -40+7y=16 +40 +40 7y=56 /7 /7 y=8

OpenStudy (luigi0210):

Plug in x for the equation: -5(8)+7y=16 -40+7y=16 7y=56 y=8

OpenStudy (anonymous):

thanks